# LV-MAE: Learning Long Video Representations through Masked-Embedding Autoencoders

### v1

@@ 7 Implementation Details
[…]
@@ 10 Additional Interpretable Predictions Results
In Fig. 5, we attach additional examples of the interpretable predictions experiment from Sec. 4.4.
@@ 11 Additional Related Work
Numerous models have been proposed for short-video understanding, achieving remarkable performance on tasks such as action recognition [2], video classification [11], text-video retrieval [33], video captioning [49], and video question answering [25]. Multimodal models like CLIP [36], InternVideo [43], and LanguageBind [54] have demonstrated strong capabilities in aligning video and language representations. In this work we leverage the output emebedding of these models to learng long-video representations.
@@
Long-video language understadning methods.

### v2

@@ 7 Implementation Details
$\text{LV-MAE}_{\text{InternVideo2}}$(LP) Semantic 55.14 58.56 49.40 43.26 72.84 37.77 51.22 52.60
$\text{LV-MAE}_{\text{InternVideo2}}$(LP) Random 57.01 57.36 52.38 49.65 70.37 39.89 48.78 53.63
$\text{LV-MAE}_{\text{InternVideo2}}$(AP) Random 54.21 66.95 55.36 56.03 80.25 42.55 53.66 58.43
$\text{LV-MAE}_{\text{LanguageBind}}$(LP) Semantic 71.03 67.47 54.76 53.90 67.90 42.02 56.09 59.02
$\text{LV-MAE}_{\text{LanguageBind}}$(LP) Random 71.96 67.12 55.36 53.19 71.60 37.76 51.22 58.32
$\text{LV-MAE}_{\text{LanguageBind}}$(AP) Random 78.50 69.17 60.12 61.70 72.84 39.36 56.09 62.54
@@ 9 Extended LVU ablation
Impact of Segment Length.
In the main paper, we report results obtained by partitioning each long video into five second clips. To assess the sensitivity of our method to this design choice, we also trained models on longer fixed-length clips (10 seconds and 15 seconds) as well as on variable-length, shot-based segments. Table LABEL:tab:segment_length summarizes the average performance across all LVU downstream tasks. The five second configuration remains the most effective, achieving an average score of 63.40 and consistently outperforming the longer and shot-based alternatives. Finally, we examined the use of overlapping five second clips; this introduces redundant context and reduces the average score by 4.5, from 63.40 to 58.90.
Table 9: Impact of Segment Length. The average Top-1 accuracy results obtained by LV-MAE on the LVU benchmark using different segment lengths.
5 seconds 10 seconds 15 seconds shots
63.40 61.43 61.05 62.17
Architecture and Clip-Length Ablations.
[…]
Table 10: Results on the LVU benchmark using different architectures and clip-length ablations.
Method FB Metadata $\uparrow$ Content $\uparrow$ Avg.
Dir. Genre Writer Year Scene Speak Rel.
ViS4mer 62.61 54.71 48.80 44.75 67.44 40.79 57.14 53.7
VideoMamba 67.29 65.24 52.98 48.23 70.37 40.43 62.50 58.1
LanguageBind-Transformer 24.30 57.88 16.07 18.44 35.80 32.45 51.22 33.7
LanguageBind-Mamba 61.68 70.38 51.78 56.74 74.04 38.83 43.90 56.8
$\text{LV-MAE}({\text{frame-MAE feature extractor)}}$ 64.49 62.50 49.4 48.23 67.90 36.70 51.22 54.3
$\text{LV-MAE}_{\text{LanguageBind}}$(Ours) 77.57 71.57 64.28 58.15 72.84 40.95 58.53 63.4
Impact of Input Frame Count.
[…]
Figure 5: Average performance of LVU benchmark rises monotonically as the frame count increases.
10 Short-Video Performance
Our approach preserves performance on short-video understanding tasks. Specifically, we preserve the accuracy of LanguageBind on Kinetics-400 (600) [22]. We achieved 78.2% on K400 and 79.1% on K600 compared to 77.6% and 79.5% in LanguageBind.
@@ 11 Additional Interpretable Predictions Results
In Fig. 6, we attach additional examples of the interpretable predictions experiment from Sec. $4.$
@@ 12 Additional Related Work
Numerous models have been proposed for short-video understanding, achieving remarkable performance on tasks such as action recognition [2], video classification [11], text-video retrieval [33], video captioning [50], and video question answering [25]. Multimodal models like CLIP [36], InternVideo [44], and LanguageBind [55] have demonstrated strong capabilities in aligning video and language representations. In this work, we leverage the output embedding of these models to learn long-video representations.
@@
Long-video language understanding methods.